OpenStudy (anonymous):

hydrogen-filled balloon was ignited and 1.50 grams of hydrogen reacted with 12.0 grams of oxygen how many grams of water vapor formed (assume that water vapor is the only product)

OpenStudy (jebonna):

There is a very easy way to do this. Due to the Law of Conservation of Mass, mass cannot be created or destroyed. This means that the mass of the reactants (in this case hydrogen and oxygen) will always equal the mass of the products (which is in this case water vapor). This means that basically, the mass of water vapor that is created is the mass of hydrogen and oxygen combined (1.5 + 12 = 13.5). So the mass would be 13.5g of water vapor. However if you would like to do these sort of equations anyway (e.g. if you only got the mass of one reactant) you can use what is written below: First off write the equation for this reaction: H2 (g) + O2(g) = H2O (g) Balance it: 2H2 (g) + O2(g) = 2H2O (g) Now, you have to work out the moles of either oxygen or hydrogen to find the moles of water vapor, to find the mass. So lets use hydrogen. To work out the moles of hydrogen you will have to do: moles = mass / molar mass Plug in the values, that will be: 1.50g / 2g mol^-1 = 0.75 mol of hydrogen Now, there are TWO lots of Hydrgogen, and TWO lots of H2O. Moles cannot be added or made, therefore the moles of hydrogen will be the same as the moles of H2O. As there are two lots of both of them (they both have a no. 2 in front of them), we do not have to change the moles at all. If we found the moles of O2, then we would have to times the moles of O2 by 2 as there is only ONE lot of O2, however there are TWO lots of H2O, and remember the moles will be the same throughout the equation. So, the moles of H2O is: 0.75 mol Now we can work out the mass of H2O using this equation: mass = moles x molar mass Plug in the values: 0.75mol x 18g mol^-1 = 13.5g So, there will be 13.5g of water vapor made. This means that the Law of Conservation of Mass is correct, and the mass of the products will equal the mass of the reactants. I hop this helps :)